What is the spin multiplicity of nitrogen in the ground state?? Give reason for your reason

Grade:12

6 Answers

manish nayak
12 years ago

for calculating the total spin of any element we have to use the formula 2[S] +1 

where [S] represents the total spin which is calculated as total no. of unpaired electrons * +/- ( 1/2  )

as ground state of N contains 3 unpaired electrons {1s22s23p3}.....

the ans is 

7

Priyavrat Sharma
11 years ago

Spin Multiplicity of nitrogen atom is 4

Purvang parmar
6 years ago
Spin multiplicity of any atom is half of its unpaired electrons so spin multiplicity of nitrogen is 1.5 as it has 3 unpaired electrons
Utkarsh
6 years ago
 
Spin multiplicity of any atom is (n+1) where (n) is the number of unpaired electrons. the no of unpaired electron in nitrogen is 3 (1s22s22p3)
n=3
(n+1)=(3+1)=4
joharzeb
6 years ago
Formula for finding spin multipicity= 2S+1 (i)
here for nitrogen electronic configuration=1s22s22p3 
according to hunds rule the unpaired electrons is 3,
So S=1/2+1/2+1/2=3/2, use this value in equation (i) so the answer is 4.
Vinoy Kumar
5 years ago
The no. Of unpaired electrons in nitrogen is 3. So total spin = +3/2 or -3/2=s.Now the spin multiplicity =2|s|+1=2×3/2+1=4 is the answer.